OpenStudy (anonymous):

It takes Maria 11 mins to weed a garden. Mike can weed the same garden in 10 minutes. How long would it take them if they work together? - I got 21 mins but i honestly don't know how to round it.

OpenStudy (saifoo.khan):

\[\frac 1x =\frac{1}{10}+\frac{1}{11}\]

OpenStudy (saifoo.khan):

1/10 = Mike's part. 1/11 = Maria'a part. 1/x = time it will take if both will work.

OpenStudy (anonymous):

Alright and so then i would find the common denominator and then solve for the x?

OpenStudy (hba):

Now take the lcm,cross multiply and get to x.

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

I got 120/110. do i divide 110 by 120 and turn it into a decimal?

OpenStudy (saifoo.khan):

Nope/

OpenStudy (anonymous):

-.- okay so for my lcd i did 11 by 10. then okay i messed up my adding. it should be 21/110

OpenStudy (saifoo.khan):

1/x = 21/110

OpenStudy (anonymous):

then i would multiply the 1 and that would isolate the x. leaving me with the 21/110 and then i simplify it?

OpenStudy (saifoo.khan):

x is in the denominator currently. Do something to bring it to the numerator.

OpenStudy (anonymous):

multiply the 110 to the top and bottom. I would get 110 over 110x?

OpenStudy (hba):

See this, \[\frac{ 1 }{ a }=\frac{ 11 }{ 10 }\] \[a=\frac{ 10 }{ 11 }\]

OpenStudy (saifoo.khan):